OpenStudy (andertj21):

$4,000 is invested at 3% interest. How much money must be invested at 5% interest so that the total interest from the two investments is $195 after a year? $1,250 $1,500 $1,750

OpenStudy (anonymous):

Okay, 4000 * 0.03 = a a + x(0.05) = 4000 + 195

OpenStudy (andertj21):

So its B?